I would like help, I am unsure how to solve the following questions from my study guide. I am no confused because I have to use an excel sheet upon solving them

1. The current cost function for a lab that evaluates blood tests is C=400,000 + 20Q, where Q is the number of tests performed annually. If the lab expects to perform 50,000 tests annually, what are the average costs per evaluation?

2. A clinic finds that it can reduce costs by eliminating appointments. The clinic is able to eliminate some of the administrative staff, and physicians become more productive. Patients simply wait until the physician is available, so the physician has virtually no down time. Should a single payer system adopt this technique? In five sentences or fewer, provide your opinion on the topic.

3. Consider the following screening for cancer applied to 100,000 people, of whom 1,000 have cancer. Each screening costs $10, picks up 85% of cancer cases, and 10% of the time falsely diagnoses cancer. Detection of cancer (rightly or wrongly) leads to a further exact diagnostic test that costs $100. Correct early detection of cancer by the test is valued at $10,000. (a) Is the first screening worthwhile? Explain your answer. (b) Is a second screening, performed on the individuals who did not get the full diagnostic test after the first screening, worthwhile? Assume that the cancer screening applied to the remaining group costs $10 a person, and the follow up confirmatory test still costs $100. Explain your answer.

4. A hospital just received a $4 million gift from a benefactor and must choose how to best spend the money. The first project involves updating the operating room equipment. It is expected that these updates would provide $600,000 in benefits at the end of each year. The new equipment has an expected lifespan of 10 years. The second project involves building a small addition to the hospital that will add 10 new patient rooms. This project is expected to provide $250,000 in benefits at the end of each year over an expected lifespan of 30 years. Using a discount rate of 5%, which project offers a larger net benefit?
Answer #1

1. Cost function : C = 400,000 + 20Q

Substituting the Value of Q as 50000, we get

C = 400,000 + 20*50000

= 400,000 + 10,00,000

= 1400000

So the Average cost per evaluation = Total Cost / No. of tests

= 1400000 / 50000

= 28

2. Single payer system covers the costs of all essential healthcare services. Eliminating appointments would create certain problems like:

  • It will create frustration amongst the target patients as they are unable to get the appointmnets. Soon they will switch to other players.
  • It will increase the waiting time for the patients and thus make restless.
  • Administrative staff also takes care of preliminary screening of the patients and measure their vital statistics. Reducing the administrative staff would increase their workload and thus patients need to wait for more hours. It would also increase the workload of the physicians.
  • Administrative staff would also be frustrated because of increased workload.
  • All this frustration would be poured on to the patients and the quality of services offered would get affected and the patients remain unsatisfied.